At normal pressure (101325 Pa) a gas has 1 ∙ 1022 molecules per unit volume. What is the average kinetic energy of its molecules?
1
Expert's answer
2021-03-02T18:05:24-0500

We can find the average kinetic energy of the molecules of a gas at STP (101325 Pa and 293.15 K) as follows:


"KE=\dfrac{3}{2}Nk_BT,""KE=\dfrac{3}{2}\cdot1.0\cdot10^{22}\ m^{-3}\cdot1.38\cdot10^{-23}\ \dfrac{J}{K}\cdot293.15\ K=60.7\ J."